# InsightBuddy-AI: Medication Extraction and Entity Linking using Large Language Models and Ensemble Learning

### v1

@@ Abstract
Medication Extraction and Mining play an important role in healthcare NLP research due to its practical applications in hospital settings, such as their mapping into standard clinical knowledge bases (SNOMED-CT, BNF, etc.). In this work, we investigate state-of-the-art LLMs in text mining tasks on medications and their related attributes such as dosage, route, strength, and adverse effects. In addition, we explore different ensemble learning methods (Stack-Ensemble and Voting-Ensemble) to augment the model performances from individual LLMs. Our ensemble learning result demonstrated better performances than individually fine-tuned base models BERT, RoBERTa, RoBERTa-L, BioBERT, BioClinicalBERT, BioMedRoBERTa, ClinicalBERT, and PubMedBERT across general and specific domains. Finally, we build up an entity linking function to map extracted medical terminologies into the SNOMED-CT codes and the British National Formulary (BNF) codes, which are further mapped to the Dictionary of Medicines and Devices (dm+d), and ICD. Our model’s toolkit and desktop applications are publicly available at https://github.com/HECTA-UoM/ensemble-NER.
@@
Pablo Romero Manchester Metropolitan University Oxford Rd Greater Manchester, UK pablo2004romero@gmail.com                        Lifeng Han and Goran Nenadic University of Manchester Oxford Rd Greater Manchester, UK first.last@manchester.ac.uk corresponding author
@@ 1 Introduction
Information Extraction on Medications and their related attributes plays an important role in natural language processing (NLP) applications in the clinical domain to support digital healthcare. Clinicians and healthcare professionals have been doing manual clinical coding for quite a long time to map clinical events such as diseases, drugs, and treatments into the existing terminology knowledge base, for instance, ICD and SNOMED. The procedure can be time-consuming yet without a guarantee of total correctness due to human-introduced errors. With the process of automated information extraction on medications, it will be further possible to automatically map the extracted terms into the current terminology database, i.e. the automated clinical coding. Due to the promising future of this procedure, different NLP models have been deployed in medication mining and clinical coding in recent years. However, they are often studied separately. In this work, we investigate text mining of medications and their related attributes (dosage, route, strength, adverse effect, frequency, duration, form, and reason) together with automated clinical coding into one pipeline. In addition, we investigate the ensemble learning mechanisms (Stack and Voting) on a broad range of NLP models fine-tuned for named entity recognition (NER) tasks. These models include both general domain trained BERT, RoBERTa, RoBERTa-L, and domain-specific trained BioBERT, BioClinicalBERT, BioMedRoBERTa, ClinicalBERT, and PubMedBERT. In this way, users do not have to worry about which models to choose for clinical NER. Instead, they can just place the newer models into the ensemble-learning framework to test their performances. We offer desktop applications and web interfaces for the clinical NER, ensemble, and coding models we are developing.
@@ 2 Literature Review and Related Work
2.1 Clinical NER and Ensemble Learning
Named Entity Recognition (NER) is a critical task for extracting key information from unstructured text, like medical letters. The complexity and context-dependency of medical language pose significant challenges for accurate entity extraction. Traditional approaches to NER, such as rule-based systems, have shown limited success in capturing the nuanced contextual information crucial for clinical NER Nadeau and Sekine (2007).
The advent of deep learning methods, particularly Long Short-Term Memory (LSTM) networks, marked a significant improvement in NER performance Graves and Schmidhuber (2005). LSTMs demonstrated an ability to capture long-range dependencies in text. However, these models still struggled with rare entities and complex contextual relationships in clinical notes.
The introduction of BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2019) in 2018 revolutionised various Natural Language Processing (NLP) tasks, including NER. BERT’s self-attention mechanism and bidirectional training allow it to capture nuanced contextual information over long pieces of text. The model’s pre-training on a large corpus using a masked language modelling objective builds rich token representations. The model can then be later fine-tuned by adding a classification layer at the end of the network to make decisions over each individual token embedding.
@@ 2.1 Clinical NER and Ensemble Learning
Despite these improvements, single-model approaches still struggle with the inherent complexity and variability of clinical text, as the comparative studies reported in Belkadi et al. (2023) across different models using BERT, ClinicalBERT, BioBERT, and scratch-learned Transformers. Ensemble methods have emerged as a promising direction to address these challenges, they have proven useful in other fields, such as computer vision Lee et al. (2018). By combining multiple models, ensembles can leverage the strengths of different models while mitigating their individual weaknesses. In the context of NER, ensembling has shown performance improvements, as shown by Naderi et al. (2021), where an ensemble is used on a health and life science corpora for a significant improvement in performance over single models. Naderi et al. (2021) conducted max voting for word-level biology, chemistry, and medicine data. However, on clinical/medical NER, they only focused on French using the DEFT benchmark dataset; while for the other two domains of biology and chemistry, they tested on English data.
There are two commonly used ensemble methods, voting and stacked ensembles.
  • Maximum voting in ensembles where each model contributes equally to the final decision as used in the paper Naderi et al. (2021) have proved effective.
  • We want to explore if training a network on the outputs of the ensemble enables it to capture more nuanced relationships. This is accomplished using a method called stacking introduced by Wolpert (1992). Stacking offers a more sophisticated approach by training a meta-model on the outputs of the base ensemble; the model is expected to learn more complex patterns from the outputs of the ensemble, leading to better predictions. This has proven effective in this paper Saleh et al. (2022) where they use a stacked ensemble with a support vector machine for sentiment analysis. We use a simple feed-forward network from the outputs of the ensemble to the final labels. more examples on stacked ensemble can be found at Mohammed and Kora (2022); Güneş et al. (2017).
While ensemble methods have shown promise in other NER tasks, their effectiveness for clinical NER, particularly on challenging datasets like n2c2 2018 Henry et al. (2020), remains unexplored. The n2c2 2018 dataset presents unique challenges due to its diverse entity types and the high degree of domain-specific knowledge required.
@@
This work aims to address this gap by investigating whether stacked and voting ensembles can improve NER performance on clinical notes, potentially overcoming the limitations of current single-model approaches.
However, it’s important to acknowledge the limitations of ensemble methods, particularly in clinical settings. The increased computational resources required for training and inference present challenges in resource-constrained environments. The trade-off between improved performance and increased computational cost must be carefully evaluated in the context of clinical workflows and available infrastructure.
2.2 Clinical Entity Linking
Entity Linking in the clinical domain is often conducted as Clinical Coding, which traditionally requires human coders to map clinical events into dictionaries or knowledge bases such as ICD, SNOMED, and BNF. However, due to the time-consuming nature of such tasks, automated coding has been explored using different methods including rule-based ones such as Farkas and Szarvas (2008) and machine learning methods such as using neural networks CNN by Mullenbach et al. (2018), Transformers by Liu et al. (2022). There are also comparative studies of different methods with explanability Glen et al. (2024). In our work, to make the extracted clinical events more valuable, we will integrate a clinical coding function into our platform by mapping into the BNF coding list, which is further mapped to ICD, SNOMED, and dm+d based on existing work.
@@ 2.2 Clinical Entity Linking
Figure 1: MedNER Framework Pipeline: individual NER model fine-tuning, ensemble, and entity linking. Two kinds of base models include the general domain and the biomedical domain with their Huggingface repositories in Figure 2. Pre-preprocessing data: cut the sequence with the first full stop “.” after the 100th token, otherwise, cut the sequence up to 128 tokens. Fine-tuning: using the same parameter sets for all eight models. Ensemble: different strategies will be displayed in Fig 3. Entity Linking: links to clinical KB including BNF and SNOMED.
Figure 2: MedNER integrated individual models and their Huggingface repositories.
@@ 3 Methodologies
The Overall framework of MedNER is shown in Figure 1, which displays the base models we included from the general domain 1) BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and RoBERTa-Large, and 2) biomedical/clinical domains BioBERT Lee et al. (2019), BioClinicalBERT Alsentzer et al. (2019), BioMedRoBERTa Gururangan et al. (2020), ClinicalBERT Wang et al. (2023), and PubMedBERT Gu et al. (2020). The fine-tuning of eight models uses the same set of parameters (Section B for parameter selections) and the n2c2-2018 shared task training data with data pre-processing. The initial evaluation phase using n2c2-2018 testing set gives an overall idea of each model’s performance. This is followed by ensemble learning on all the models’ outputs. With the output from NER models, we add an entity linking function to map the extracted medical entities into the standard clinical terminology knowledge base (KB), using SNOMED-CT and BNF as our initial KB. Finally, we also offer user-friendly desktop applications and interfaces for the MedNER project.
For data pre-processing, we chunk the sequence into a maximum of 128 tokens. If there is a full stop “.” between the 100th and 128th tokens, it will be cut at the full stop. Regarding ensemble-learning strategy, we draw a MedNER Ensemble figure (Figure 3) to explain in detail. Firstly the initial output of eight individual fine-tuned NER models is tokenised, i.e. at the sub-word level, due to the model learning strategy, e.g. “Para ##ce ##tam ##ol” instead of “Paracetamol”. What we need to do at the first step is to group the sub-word tokens into words for both practical application and voting purposes. However, each sub-word is labeled with predefined labels and these labels often do not agree with each other within the same words. We designed three group solutions, i.e. first-token voting/selection, max-token voting, and average voting. The first-token voting is to assign a word the same label as its first sub-word piece. For example, using this strategy, the word “Paracetamol” will be labeled as “B-Drug” if its first sub-word “Para” is labeled as “B-Drug” regardless of other labels from the subsequent sub-words. The max-token voting will assign a word the label that has the highest sub-word logit, this indicates that the model is more confident in that prediction, the higher the logit is. The average voting solution calculates the average logits across all sub-words predictions and then samples from this to get the label for the entire word.
@@
Figure 3: MedNER Ensemble Pipeline: individual NER model fine-tuning outputs are at token/sub-word level. "Logits are the outputs of a neural network before the activation function is applied" first, we do the grouping of sub-words into words using three strategies: first token label, max token voting, or average voting (from our results, the first-token-lable selection gives higher Recall, while other two voting give higher precision, but they all end with the same F1 score, ref Figure xxx ). $|$ We take the best output from the first token label selection as the solution. For word-level ensemble on eight models, we have two solutions for voting, 1) either majority voting with >= 4 labels as the same then we pick it, otherwise choose default “O”, or 2) max voting with the most popular label whatever it is; for max voting, if there is a tie, e.g. (3,3,2), we tested both alphabetical pick-up, or random pick-up of tied labels. Our results show that “>=4, or O” performs similarly to “max + alphabetical”, while “max + random” slightly performs lower.
We also draw the Stacked-Ensemble in Figure 4 and 5, where the model training and one-hot encoding are illustrated. In the training phase, we slit the real data into 80% and 20% for model training and testing. Model exports are conducted only if at least 2 models are predicting a label that is not “O”; otherwise “O” is the default option. For model deployment, output logis for each model are turned into a one-hot encoded vector, concatenated and saved along with the real label for each token. There are 8 one-hot encoded vectors from 8 individual models and 1 label. To use one-hot encoding instead of the logit itself is to avoid the model overfitting.
Figure 4: StackEnsemble: training strategy.
Figure 5: StackEnsemble: one-hot encoding data. Instead of feeding the first logit, which produced a lower score due to overfitting, we now feed the one-hot encoding, which produced the highest Precision.
@@ 4 Experimental Evaluations
We use the n2c2-2018 shared task data on NER of adverse drug events and related medical attributes Henry et al. (2020). The data is labels with the following list of labels: ADE, Dosage, Drug, Duration, Form, Frequency, Reason, Route, and Strength in BIO format. So overall we have 19 labels, 2 (B/I) x 9 + 1 (O). The original training and testing sets are 303 and 202 letters respectively. We divided the original training set into two parts (9:1 ratio) for our model selection purposes: our new training and validation set, following the data split from recent work by Belkadi et al. (2023).
@@
We report Precision, Recall, and F1 score in two categories “macro” and “weighted”, in addition to Accuracy. The “macro” category treats each label class the same weight regardless of their occurrence rates, while the “weighted” category” assigns each label class with a weight according to their occurrence in the data. We first report the individual model fine-tuning scores and compare them with related work (sub-word level); then we report the ensemble model evaluation with different ensemble solutions (word-level).
Figure 6: ClinicalBERT Eval at Sub-word Level. This score is similar to Belkadi et al. (2023) paper on ClinicalBERT-Apt whose macro: .842, .834, .837; and weighted: .974, .974, .974) which says our fine-tuning is successful. However, our best domain-specific model BioMedRoBERTa produces higher score: macro P/R/F (0.8482 0.8477 0.8468) and weighted P/R/F (0.9782 0.9775 0.9776) and Accuracy 0.9775 as in Figure 8. Furthermore, the fine-tuned RoBERTa-L even achieved higher scores of (0.8489 0.8606 0.8538) for P/R/F1 and Acc 0.9782 in Figure 17. Afterwards, in this paper, we emphasis on word level instead of sub-word, which was focused by Belkadi et al. (2023).
Figure 7: MedNER individual sub-word level model eval on n2c2-2018 test set.
4.1 Individual Models Finetuned: sub-word level
@@
The performances of individual models after fine-tuning are reported in Figure 7, where it says that RoBERTa-L performs that best on macro Precision (0.8489), Recall (0.8606), and F1 (0.8538) score across general domain models, also winning domain-specific models. BioiMedRoBERTa wins the domain-specific category models producing macro Precision, Recall, and F1 scores (0.8482 0.8477 0.8468). In comparison to the NER work from Belkadi et al. (2023), who’s macro avg scores are: 0.842, 0.834, 0.837 from ClinicalBERT-Apt, our fine-tuned ClinicalBERT has similar / comparable performances (0.848, 0.825, 0.834), which shows our fine-tuning was successful. However, our best domain-specific model BioMedRoBERTa produces higher scores: macro P/R/F1 (0.8482 0.8477 0.8468) and weighted P/R/F1 (0.9782 0.9775 0.9776) and Accuracy 0.9775 as in Figure 6. Furthermore, the fine-tuned RoBERTa-L even achieved higher scores of (0.8489 0.8606 0.8538) for macro P/R/F1 and Acc 0.9782 in Figure 13. Both fine-tuned BioMedRoBERTa and RoBERTa-Large also win the best models reported by Belkadi et al. (2023) which is their ClinicalBERT-CRF model, macro avg (0.85, 0.829, 0.837), Acc 0.976. Afterwards, in this paper, we emphasis on word level instead of sub-word, which was focused on by Belkadi et al. (2023).

### v2

@@ Abstract
Medication Extraction and Mining play an important role in healthcare NLP research due to its practical applications in hospital settings, such as their mapping into standard clinical knowledge bases (SNOMED-CT, BNF, etc.). In this work, we investigate state-of-the-art LLMs in text mining tasks on medications and their related attributes such as dosage, route, strength, and adverse effects. In addition, we explore different ensemble learning methods (Stack-Ensemble and Voting-Ensemble) to augment the model performances from individual LLMs. Our ensemble learning result demonstrated better performances than individually fine-tuned base models BERT, RoBERTa, RoBERTa-L, BioBERT, BioClinicalBERT, BioMedRoBERTa, ClinicalBERT, and PubMedBERT across general and specific domains. Finally, we build up an entity linking function to map extracted medical terminologies into the SNOMED-CT codes and the British National Formulary (BNF) codes, which are further mapped to the Dictionary of Medicines and Devices (dm+d), and ICD.
@@
Pablo Romero Manchester Metropolitan University Oxford Rd Greater Manchester, UK pablo2004romero@gmail.com                        Lifeng Han LIACS & LUMC Leiden University, NL University of Manchester, UK l.han@lumc.nl corresponding author Software Download Link                        Goran Nenadic University of Manchester Oxford Rd Greater Manchester, UK g.nenadic@manchester.ac.uk
@@ 1 Introduction
Information Extraction on Medications and their related attributes plays an important role in natural language processing (NLP) applications in the clinical domain to support digital healthcare. Clinicians and healthcare professionals have been doing manual clinical coding for quite a long time to map clinical events such as diseases, drugs, and treatments into the existing terminology knowledge base, for instance, ICD and SNOMED. The procedure can be time-consuming yet without a guarantee of total correctness due to human-introduced errors. With the process of automated information extraction on medications, it will be further possible to automatically map the extracted terms into the current terminology database, i.e. the automated clinical coding. Due to the promising future of this procedure, different NLP models have been deployed in medication mining and clinical coding in recent years. However, they are often studied separately. In this work, 1) we investigate text mining of medications and their related attributes (dosage, route, strength, adverse effect, frequency, duration, form, and reason) together with automated clinical coding into one pipeline. In addition, 2) we investigate the ensemble learning mechanisms (Stack and Voting) on a broad range of NLP models fine-tuned for named entity recognition (NER) tasks. These models include both general domain trained BERT, RoBERTa, RoBERTa-L, and domain-specific trained BioBERT, BioClinicalBERT, BioMedRoBERTa, ClinicalBERT, and PubMedBERT. In this way, users do not have to worry about which models to choose for clinical NER. Instead, they can just place the newer models into the ensemble-learning framework to test their performances.
@@ 2 Literature Review and Related Work
Named Entity Recognition (NER) is a critical task for extracting key information from unstructured text, like medical letters. The complexity and context-dependency of medical language pose significant challenges for accurate entity extraction. Traditional approaches to NER, such as rule-based systems, have shown limited success in capturing the nuanced contextual information crucial for clinical NER Nadeau and Sekine (2007). The advent of deep learning methods, particularly Long Short-Term Memory (LSTM) networks, marked a significant improvement in NER performance Graves and Schmidhuber (2005), e.g. the ability to capture long-range dependencies in text. However, these models still struggled with rare entities and complex contextual relationships in clinical notes. The introduction of BERT (Bidirectional Encoder Representations from Transformers) Devlin et al. (2019) revolutionised various NLP tasks, including NER. The model’s pre-training on a large corpus using a masked language modelling objective builds rich token representations. The model can then be later fine-tuned by adding a classification layer at the end of the network to make decisions over each individual token embedding.
@@
Despite these improvements, single-model approaches still struggle with the inherent complexity and variability of clinical text, as the comparative studies reported in Belkadi et al. (2023) across different models using BERT, ClinicalBERT, BioBERT, and scratch-learned Transformers. Ensemble methods have emerged as a promising direction to address these challenges, they have proven useful in other fields, such as computer vision Lee et al. (2018). By combining multiple models, ensembles can leverage the strengths of different models while mitigating their individual weaknesses. In the context of NER, ensembling has shown performance improvements, as shown by Naderi et al. (2021), where an ensemble is used on a health and life science corpora for a significant improvement in performance over single models. Naderi et al. (2021) conducted max voting for word-level biology, chemistry, and medicine data. However, on clinical/medical NER, they only focused on French using the DEFT benchmark dataset; while for the other two domains of biology and chemistry, they tested on English data. There are two commonly used ensemble methods, voting and stacked ensembles: 1) Maximum voting in ensembles where each model contributes equally to the final decision as used in the paper Naderi et al. (2021) have proved effective. This is where the most voted label is picked. 2) Training a network on the outputs of the ensemble aims to capture more nuanced relationships. This is accomplished using a method called stacking introduced by Wolpert (1992). Stacking offers a more sophisticated approach by training a meta-model on the outputs of the base ensemble; the model is expected to learn more complex patterns from the ensemble outputs, leading to better predictions. This has proven effective in this paper Saleh et al. (2022) where they use a stacked ensemble with a support vector machine (SVM) for sentiment analysis. Instead, we will use a simple feed-forward network from the outputs of the ensemble to the final labels for our tasks. more examples on stacked ensemble can be found at Mohammed and Kora (2022); Güneş et al. (2017).
@@
While ensemble methods have shown promise in other NER tasks, their effectiveness for clinical NER, particularly on challenging datasets like n2c2 2018 Henry et al. (2020), remains unexplored. This work aims to address this gap by investigating whether stacked and voting ensembles can make an impact on NER tasks on clinical notes, potentially overcoming the limitations of current single-model approaches.
@@
Figure 1: InsightBuddy Framework Pipeline: individual NER model fine-tuning, ensemble, and entity linking. Two kinds of base models include the general domain and the biomedical domain with their Huggingface repositories in Table LABEL:tab:EnsembleNER-model-list. Pre-preprocessing data: cut the sequence with the first full stop “.” after the 100th token, otherwise, cut the sequence up to 128 tokens. Fine-tuning: using the same parameter sets for all eight models. Ensemble: different strategies will be displayed in Fig 7. Entity Linking: links to clinical KB including BNF and SNOMED.
@@ 3 Methodologies
The Overall framework of InsightBuddy is shown in Figure 1, which displays the base models we included from the general domain 1) BERT Devlin et al. (2019), RoBERTa Liu et al. (2019), and RoBERTa-Large, and 2) biomedical/clinical domains BioBERT Lee et al. (2019), BioClinicalBERT Alsentzer et al. (2019), BioMedRoBERTa Gururangan et al. (2020), ClinicalBERT Wang et al. (2023), and PubMedBERT Gu et al. (2020). The fine-tuning of eight models uses the same set of parameters (Section D for parameter selections) and the n2c2-2018 shared task training data with data pre-processing. The initial evaluation phase using n2c2-2018 testing set gives an overall idea of each model’s performance. This is followed by ensemble learning on all the models’ outputs. With the output from NER models, we add an entity linking function to map the extracted medical entities into the standard clinical terminology knowledge base (KB), using SNOMED-CT and BNF as our initial KB, which is further mapped to ICD and dm+d.
For data pre-processing, we chunk the sequence into a maximum of 128 tokens. If there is a full stop “.” between the 100th and 128th tokens, it will be cut at the full stop. Regarding ensemble-learning strategy, we draw a InsightBuddy Ensemble figure (Figure 7) to explain in detail. Firstly the initial output of eight individual fine-tuned NER models is tokenised, i.e. at the sub-word level, due to the model learning strategy, e.g. “Para ##ce ##tam ##ol” instead of “Paracetamol”. What we need to do at the first step is to group the sub-word tokens into words for both practical application and voting purposes. However, each sub-word is labeled with predefined labels and these labels often do not agree with each other within the same words. We designed three group solutions, i.e. first-token voting/selection, max-token voting, and average voting. The first-token voting is to assign a word the same label as its first sub-word piece. For example, using this strategy, the word “Paracetamol” will be labeled as “B-Drug” if its first sub-word “Para” is labeled as “B-Drug” regardless of other labels from the subsequent sub-words. The max-token voting will assign a word the label that has the highest sub-word logit, this indicates that the model is more confident in that prediction, the higher the logit is. The average voting solution calculates the average logits across all sub-words predictions and then samples from this to get the label for the entire word.
@@
We also draw the Stacked-Ensemble in Figure 8 and 9, where the model training and one-hot encoded model predictions are illustrated. In the training phase, we cut the real data into 80% and 20% for the training and testing of the model. Model exports are conducted only if at least 2 models are predicting a label that is not “O”; otherwise “O” is the default option and the output is ignored and not included in the stacked training data. For training data collection, output logits for each model are converted into a one-hot encoded vector, concatenated and saved along with the real label for each token. There are 8 one-hot encoded vectors from 8 individual models and 1 label. So the model during training will see the value “1” eight times from the eight models, and the value “0” for the rest of the vector values. Overall, there are 8 vectors with each length of 19 digits. So there will be 8 (number of models) $\times$× 19 (number of labels) - 8 (eight 1s as there are 8 one hot encoded vectors so they have a single 1 each) = 144 “0” values for every training example. We use one-hot encoding instead of the output logits themselves to avoid the model overfitting because the model makes more confident predictions when running on the training set. As this is the data that it was originally trained on, it is very confident with it’s predictions. We can mitigate this by only feeding the one-hot encoded vectors to the stacked network.
@@ 4 Experimental Evaluations
We use the n2c2-2018 shared task data on NER of adverse drug events and related medical attributes Henry et al. (2020). The data is labeled with the following list of labels: ADE, Dosage, Drug, Duration, Form, Frequency, Reason, Route, and Strength in BIO format. So, overall, we have 19 labels, 2 (B/I) x 9 + 1 (O). The original training and testing sets are 303 and 202 letters respectively. We divided the original training set into two parts (9:1 ratio) for our model selection purposes: our new training and validation set, following the data split from recent work by Belkadi et al. (2023).
@@
We report Precision, Recall, and F1 score in two categories “macro” and “weighted”, in addition to Accuracy. The “macro” category treats each label class the same weight regardless of their occurrence rates, while the “weighted” category” assigns each label class with a weight according to their occurrence in the data. We first report the individual model fine-tuning scores and compare them with related work (subword level); then we report the ensemble model evaluation with different ensemble solutions (word level).
4.1 Individual Models: sub-word level
@@
The performance of individual models after fine-tuning is reported in Table 3 where it says that RoBERTa-L performs the best in the macro Precision (0.8489), Recall (0.8606) and F1 (0.8538) score across general domain models, also winning domain-specific models. BioiMedRoBERTa wins the domain-specific category models producing macro Precision, Recall, and F1 scores (0.8482 0.8477 0.8468). In comparison to the NER work from Belkadi et al. (2023), who’s macro avg scores are: 0.842, 0.834, 0.837 from ClinicalBERT-Apt, our fine-tuned ClinicalBERT has similar performances (0.848, 0.825, 0.834), which shows our fine-tuning was successful. However, our best domain-specific model BioMedRoBERTa produces higher scores: macro P/R/F1 (0.8482 0.8477 0.8468) and weighted P/R/F1 (0.9782 0.9775 0.9776) and Accuracy 0.9775 as in Figure 6. Furthermore, the fine-tuned RoBERTa-L even achieved higher scores of (0.8489 0.8606 0.8538) for macro P/R/F1 and Acc 0.9782 in Figure 13. Both fine-tuned BioMedRoBERTa and RoBERTa-Large also win the best models reported by Belkadi et al. (2023) which is their ClinicalBERT-CRF model, macro avg (0.85, 0.829, 0.837), Acc 0.976. Afterwards, in this paper, we emphasis on word level instead of sub-word, which was focused on by Belkadi et al. (2023).
4.2 Ensemble: word-level grouping (logits)
We tried first logit voting, max voting, and average voting to group sub-words into words with corresponding labels. Their results are shown in Table 1, in the upper group. First logit voting produced a higher Recall 0.8260 while Max logit voting produced a higher Precision 0.8261 resulting in higher F1 0.8232, i.e. Max logit > First logit > Average logit with macro F1 (0.8232, 0.8229, 0.8227). However, overall, their performance scores are very close, so we chose the first-logit voting output for the afterward word-level ensemble due to computational convenience.
4.3 Ensemble: Voting vs Stacked (one-hot)
Regarding Stacked Ensemble using one-hot encoded vectors, as shown in the middle group in Table 1, it actually produced higher Precision score 0.8351 in comparisons to the highest Precision 0.8261 from Voting Ensembles. However, the Recall score on macro avg is 2 point lower than the voting ensemble, 0.8065 vs 0.8260, which means that the Stacked Ensemble reduced the false positive errors but also increased the false negative error prediction. This implies that it has stricter constraint on positive predictions.
4.4 Ensemble Models: BIO-span vs non-strict word-level
So far, we have been reporting the evaluation scores on the BIO-strict label categorization, i.e. we distinguish between the label’s beginning or the inner part of the label. For instance, a B-Drug will be different from an I-Drug and it will be marked as wrong if they are different from the reference. However, we think in practice, there are situations when users do not need the BIO, especially B and I difference. In Table 1, we can see that, without considering the label difference of B and I, only focusing on the 9 label categories, word level ensemble model produced much higher Macro avg evaluations cores on Precision (0.8844) and Recall (0.8830) leading to higher F1 (0.8821), in comparison to BI-distinguished Macro F1 0.8232 (voting-max-logit) and F1 0.8156 (stacked-first-logit).
4.5 Word-level: voting ensembles vs individual fine-tuned
As in Table LABEL:tab:individual-vs-ensemble-max-voting-word-level, BioMedRoBERTa individual word level max logit grouping scores macro avg P/R/F1 (0.8065 0.8224 0.8122 563329) vs max logit ensemble voting P/R/F1 (0.8261 0.8259 0.8232), we can see that ensemble boosted P (0.8261-0.8065)/0.8065= 2.43%, and F1 (0.8232-0.8122)/0.8122= 1.35% which says the ensemble voting is successful. By increasing the Precision score, the ensembles reduce the false positive labels in the system output, while keeping the Recall at the same level, i.e. the true positive labels.
Voting Average Ensemble word level (BIO)
Metric P R F1
accuracy 0.9796
macro avg 0.8253 0.8256 0.8227
weighted avg 0.9807 0.9796 0.9798
Voting First logit Ensemble word level (BIO)
Metric P R F1
accuracy 0.9796
macro avg 0.8255 0.8260 0.8229
weighted avg 0.9807 0.9796 0.9798
Voting Max logit Ensemble word level (BIO)
Metric P R F1
accuracy 0.9796
macro avg 0.8261 0.8259 0.8232
weighted avg 0.9807 0.9796 0.9798
Stacked Ensemble first logit word level (BIO)
Metric P R F1
accuracy 0.9796
macro avg 0.8351 0.8065 0.8156
weighted avg 0.9800 0.9796 0.9794
Non-BIO-only-word ensemble
Metric P R F1
accuracy 0.9839
macro avg 0.8844 0.8830 0.8821
weighted avg 0.9840 0.9839 0.9838
Table 1: word-level grouping ensemble voting, max > first > average logit voting though they are very close scores. The stacked ensemble has the highest Precision scores, but the lowest Recall scores, which lead to lower F1. In the bottom cluster, it is the word-level evaluation without distinguishing B/I labels, evaluation on n2c2 2018 test data.
5 Entity Linking: BNF and SNOMED
To map the identified named entities into the clinical knowledge base. We use the existing code mapping sheet from the British National Formulary (BNF) web between SNOMED-CT, BNF, dm+d, and ICD  [^2] https://www.nhsbsa.nhs.uk/prescription-data/understanding-our-data/bnf-snomed-mapping. We preprocessed the SNOMED code from 377,834 to 10,804 to filter repeated examples between the mapping of SNOMED and BNF. We looked for non-drug words present in the text, then we filtered the drugs further by seeing if words like [’system’, ’ostomy’, ’bag’, ’filter’, ’piece’, ’closure’] were present in the text, and if so, it was discarded.
For SNOMED CT mapping, we applied a fuzzy search to the cleaned mapping list with drug names. Then the SNOMED CT code will be added to the searching function on the SNOMED CT web, whenever there is a match. For BNF mapping, the linking function uses keyword search to retrieve the BNF website with corresponding drugs, due to its different searching features in comparison to the SNOMED-CT web page. Potential users can select whichever is suitable to their preferences between the two clinical knowledge bases (KBs), Figure 2.
6 Discussion and Conclusion
In this paper, we investigated Stacked Ensemble and Voting Ensemble on medical named entity recognition tasks using eight pretrained LMs from both general and biomed/clinical domains. Our experiments show that our fine-tuned best individual models outperformed the state-of-the-art on standard shared task data n2c2-2018. The two ensemble strategies using output logits and one-hot encoding further improved the model performances. We also offer desktop applications and user interfaces for individual fine-tuned models where we added the entity linking/normalisation function to BNF and SNOMED CT clinical knowledge base. We call the package InsightBuddy-AI, which will be released upon acceptance.
Limitations
Ensemble methods, especially with large models, can be computationally intensive. We indeed encountered limitations regarding computational resources during training and inference. Future work includes addressing the computational overhead of ensembles, exploring other ensemble techniques, and applying the methods to different datasets.
Ethics
Appendix A InsightBuddy-AI Desktop Application
We illustrate the Desktop Applications of InsightBuddy-AI in Figure 4 and 5, for demonstration of clinical event recognition using a synthetic letter via 1) loading our pre-trained model and common NER categories via 2) loading a Huggingface NER model. There is also a sliding window feature called “context length” to allow flexible length of context around the entities visible to users, as in Figure 6. For Clinical Coding (entity linking) options, the desktop application can currently directly link the extracted entities to BNF and SNOMED-CT, as in Figure 3 from the screenshots. The InsightBuddy-AI software supports both Mac and Windows systems.
Figure 3: Choice of BNF and SNOMED-CT Linking
Figure 4: Demonstration of Clinical Events Outputs using A Synthetic Letter.
Appendix B Diagrams and Scoring Tables
B.1 Word-level vs Sub-word Level scores
From word-level ensemble result in Figure 15, it says that the ensembled model can achieve word-level evaluation scores 0.826, 0.826, and 0.823 for macro P/R/F1, which is close to sub-word level best model 0.847 F1. We can see that at word-level evaluation, there are 563,329 support tokens in Figure 15, vs sub-word level 756,014 tokens in Figure 11.
Word-level ensemble voting, max-logit voting > first-logit > average-logit, as shown in Table 1, with Macro F1 scores (0.8232, 0.8229, 0.8227) respectively, which are very close though. They have the same weighted average F1 and Accuracy scores (0.9798, 0.9796) respectively.
B.2 Ensemble: Stacked using output logits (non one-hot)
When we used the ‘output logits’ instead of ‘one-hot encoding’ for stacked ensemble, as we discussioned in the methodology section, it will lead to overfitting issues. We use the Max logit stacked ensemble as an example, in figure 17, which shows that the Stacked Ensemble using output logits produced much lower evaluation scores macro avg (0.6863 0.7339 0.6592) than the voting mechanism macro avg (0.8261 0.8259 0.8232) for (P, R, F1). The corresponding confusion matrix from the stacked ensemble using the max logit is shown in Figure 18 with more errors spread in the image, the coloured numbers outside the diagonal line.
Figure 7: InsightBuddy Voted Ensemble Pipeline: individual NER model fine-tuning outputs are at token/sub-word level. "Logits are the outputs of a neural network before the activation function is applied" first, we do the grouping of sub-words into words using three strategies: first token label, max token voting, or average voting (from our results, the first-token-lable selection gives higher Recall, while other two voting give higher precision, but they all end with the same F1 score, ref Table 1 ). $|$ We take the best output from the first token label selection as the solution. For word-level ensemble on eight models, we have two solutions for voting, 1) either majority voting with >= 4 labels as the same then we pick it, otherwise choose default “O”, or 2) max voting with the most popular label whatever it is; for max voting, if there is a tie, e.g. (3,3,2), we tested both alphabetical pick-up, or random pick-up of tied labels. Our results show that “>=4, or O” performs similarly to “max + alphabetical”, while “max + random” slightly performs lower.
Figure 8: StackedEnsemble: training strategy.
Figure 9: StackedEnsemble: one-hot encoding data. Instead of feeding the first logit, which produced a lower score due to overfitting, we now feed the one-hot encoding, which produced the highest Precision.
Figure 10: ClinicalBERT Eval at Sub-word Level. This score is similar to Belkadi et al. (2023) paper on ClinicalBERT-Apt whose macro: .842, .834, .837; and weighted: .974, .974, .974) which says our fine-tuning is successful. However, our best domain-specific model BioMedRoBERTa produces higher score: macro P/R/F (0.8482 0.8477 0.8468) and weighted P/R/F (0.9782 0.9775 0.9776) and Accuracy 0.9775 as in Figure 11. Furthermore, the fine-tuned RoBERTa-L even achieved higher scores of (0.8489 0.8606 0.8538) for P/R/F1 and Acc 0.9782 in Table LABEL:tab:individual-vs-ensemble-max-voting-word-level. Afterwards, in this paper, we emphasis on word level instead of sub-word, which was focused by Belkadi et al. (2023).